Square root of x+Under square root of x.+ under square root of x and so on = square root of x*under square root of x and do on then value of x is

Grade:9

1 Answers

Vikas TU
6 years ago
Dear Student,
√(x+√x+….. = √(x.√x…..= p 
           So, from the first one we get
           P^2= x + p
           From the second one
          P^2= x.p
          So, x= P
         Putting value we get
          P^2 = 2p
         So, the value of P is 0 or 2.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)